# Vulnerability Disclosure or Notification? Best Practices for Reaching Stakeholders at Scale

### v1

@@ 5.2.3. Messaging Infrastructure
A finder will build their messaging infrastructure or delegate it to support organisations based on the selected communication channel. The common practice of outgoing email and website is using a registered individual or disclosure-specific account with the domain name of the finder’s organisation to increase the delivery rate and trust of recipients (Li et al., 2016; Stock et al., 2016a, 2018b; Chen et al., 2024). The email account selection of each work is presented in Table 3.
@@
The dedicated account practice is to mitigate the email notification challenge of spam and phishing mitigation nowadays. Due to the increasing spam and phishing, more and more measures are in place to prevent the delivery of these messages. In the early 2000s, it was possible to spin up a mail server and immediately send out an email notification to thousands of recipients. These days, there are many different standards associated with sending out emails (SPF, DKIM, DMARC, etc.) (Li et al., 2016), and the reputation of the mail server is taken into account before an email is delivered.
@@
Although a finder or its organisation can maintain their messaging infrastructure, it is also reported in (Chen et al., 2024) that their email infrastructure is partially outsourced to a mail server, Microsoft Exchange. Their mail server introduces several restrictions, such as limited account control and email-sending rate, which prevent sending out large amounts of messages (Microsoft, [n. d.]) efficiently. This makes it harder for researchers to perform large-scale notification operations using email. This also resonates with the trends of vendors outsourcing vulnerability platforms, which limit the reachability from the sender to the responsible parties, the client of the outsourced platforms (Boucher and Anderson, 2022).
@@
There are disclosure and notification operations delegating the messaging to national CSIRTs (such as NCSC-NL), CERT/CC with VINCE (CERT/CC, [n. d.]), and Shadowserver Foundation. These are listed as Coordinator Delegation and Shadowserver Foundation in Table 3. While VINCE serves as a disclosure database with notification toward vendors (Boucher and Anderson, 2022; Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]), national CSIRTs have their mailing list or website to inform vendors and end-users (Li et al., 2016; Stock et al., 2016a; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]). In particular, Shadowserver has become a newer notification delegation option to its subscribers, including vendors and end-users (The Shadowserver Foundation, [n. d.]), such as ISPs (Çetin et al., 2019b), domain owners, and network operators (Beitis and Vanhoef, [n. d.]). Besides, if a finder works with an organisation with dedicated communication channels, such as an ISP notifying its customers, there can be a more efficient way to trigger action from the recipients (Çetin et al., 2019b; Rodríguez et al., 2021). A similar notification can also be seen in (Li et al., 2016; Çetin et al., 2017a; Chen et al., 2024), where domain name owners and cloud providers use their notification systems to prompt their customers, website or service owners, to take action.
5.2.4. Disclosure Message
@@
The formulation of the disclosure message can determine the attention of the recipients and whether or not to respond and take action (Wogalter et al., 1999). The tradeoffs of length and details of content, such as remediation and security suggestions, are extensively discussed in (Çetin et al., 2017a; Stock et al., 2016a, 2018b; Çetin et al., 2019b; Li et al., 2016; Chen et al., 2024; Stivala et al., 2024), where they share not only their disclosure message templates but also the feedback from stakeholders.
@@
The approach in  (Çetin et al., 2017a; Stock et al., 2016a, 2018b) is to provide brief vulnerability information, affected systems or hosts, and their disclosure purposes in the initial messages. The message prompted the recipients either to respond to their emails or to check a webpage for extended information on vulnerabilities and remediations with the automatically generated token of each message by the finders. This way, the sender could study the response rate and read the message with a dedicated web backend. Moreover, the finders can lower the chance of information leak in case of the wrong recipients or parties. Furthermore, disclosure messages could be embedded with HTML content, such as the logo of the disclosure party, to further check if a recipient loads the full message. However, due to the spam filter and recent email client loading mechanism, extensive embedded HTML content is no longer a suggested method confirmed in (Stock et al., 2018b; Çetin et al., 2017a). From earlier work  (Çetin et al., 2017a; Stock et al., 2018b; Li et al., 2016) to recent publications (Stivala et al., 2024; Chen et al., 2024), it has been confirmed that plain text is the suggested way with less distraction and distinction from phishing messages.
@@
Lastly, in large-scale vulnerability notifications, providing information such as domain names, IP addresses, ports, and issues found in the vulnerable systems is recommended to help stakeholders investigate issues in time, as documented from the stakeholders’ feedback in (Li et al., 2016; Stock et al., 2018b; Çetin et al., 2017a; Chen et al., 2024). However, receiving parties may have certain mail filters for incoming messages, which strictly limit the message length and attachments (Chen et al., 2024). This could hinder the case of cloud providers or domain owners with a large number of vulnerable systems running for their clients. Such stakeholders may decline the messages containing longer vulnerable host lists or attachments. This eventually makes the decision on the content more challenging for finders (Çetin et al., 2017a; Chen et al., 2024; Stivala et al., 2024).
5.2.5. Disclosure Policy Implementation
@@
Organisations present support for CVD by publishing a disclosure policy on their websites, using security.txt(Foudil and Shafranovich, 2022), or providing bounty or tracking programs (Boucher and Anderson, 2022; Moura and Heidemann, 2023; Vanhoef, [n. d.]). Boucher et al. (Boucher and Anderson, 2022) observed vendors outsourcing their programs to third-party platforms that reveal their own policies and prioritisation on vulnerability selections. They noted that the policies and preferences of the outsourced platforms may limit the incentive and direct message delivery to the responsible parties. In the vulnerability notification, it has been observed in  (Chen et al., 2024) that most informed stakeholders still do not provide security.txt or equivalent information in their responses, but privacy policies that do not necessarily help the disclosure process. The implementation of the disclosure policies among stakeholders still requires attention.
@@
Google Project Zero started with an outgoing vulnerability disclosure policy describing the timelines they would use in disclosing vulnerabilities. On the one hand, this has pressured vendors to work on mitigation and publish it within 90 days of the timeline. After several years, the 90-day deadline has become an accepted practice. Academic researchers have also started using outgoing vulnerability disclosure policies (Reidsma et al., 2023), which help build trust between stakeholders and coordinate disclosure operations. The outgoing policy has been implemented in (Van Hove et al., 2023; Chen et al., 2024), where the finders presented their procedures on vulnerability handling, notification frequency, setting public disclosure schedule, and more (National Cyber Security Centre of The Netherlands (2018), NCSC-NL; van der Ham et al., 2023) in their outgoing messages. This gives the receiving party a brief yet informative message for potential procedures and conversation. Other related work  (Çetin et al., 2017a, 2019b; Stock et al., 2018b; Moura and Heidemann, 2023; Stivala et al., 2024) did not document a dedicated disclosure policy, yet did provide equivalent information and contact information in their messages for coordinating the disclosure or exempting from the notifications.
For the finders that chose coordinator delegation, ISP intervention, and Shadowserver Foundation, they also followed and presented the disclosure policy from the support organizations (Boucher and Anderson, 2022; Çetin et al., 2019b; Van Hove et al., 2023; Moura and Heidemann, 2023; Rodríguez et al., 2021; Stock et al., 2018b, 2016a; Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]). These delegated parties often use existing relations with stakeholders and customers to create a trustworthy communication channel.
@@ 5.2.6. Post-Review
The remediation rate of the vulnerable systems is a gripe in most work, as shown in Table 2. Although there are 57% of the vulnerable hosts patched with email notification (Durumeric et al., 2014), which is relatively higher than others with remediation rates from lower than 2.25% to up to around 30% (Çetin et al., 2017a; Stock et al., 2018b, 2016a; Li et al., 2016; Chen et al., 2024; Stivala et al., 2024), it’s worth noting that the notification was 3 weeks after the notable public disclosure. Besides, the finder had to drop 56% of the detected vulnerable hosts due to responsible administrators likely having no access to treat the embedded devices (Durumeric et al., 2014). Yet, with a dedicated communication channel and trustworthy notification organisation (ISP Intervention) as in (Çetin et al., 2019b; Rodríguez et al., 2021), the remediation rate can be significantly higher; nonetheless, such a scenario requires selective recipients or extra capacity and won’t necessarily fit other end-user notifications (Çetin et al., 2019b). Moreover, not every disclosure operation can measure the remediation rate due to the nature of the vulnerabilities and affected parties (Boucher and Anderson, 2022; Van Hove et al., 2023; Moura and Heidemann, 2023; Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]). However, the operation in (Beitis and Vanhoef, [n. d.]) includes both disclosure and notification operations with a measurable impact scale. It does not provide a remediation rate since it’s still a work in progress. Their result is worth observing in the near future.
@@
As (Moura and Heidemann, 2023) states, ”Vulnerability Disclosure Considered Stressful”(title). The lack of best practices in large-scale vulnerability disclosure, and especially notification, has led to finders struggling to prepare and handle the aforementioned disclosure challenges at each stage. This practice gap has motivated the finders in the selected work to adopt existing guidelines, reflect on real operations, and contribute to best practices. However, the stress and frustration of the finders deserve attention. Whether it is disclosure or notification, various vulnerability platforms (Boucher and Anderson, 2022), tracking systems (Moura and Heidemann, 2023), ticketing systems (Li et al., 2016; Çetin et al., 2017a; Chen et al., 2024) have increased the workload of disclosure parties to deliver the messages to the responsible parties in the complex multi-party scenarios. The inaccuracy of existing abuse and generic contact information has caused false positives in contact retrieval and information leaks to the wrong recipients, which is mentioned in nearly all selected work using email as a communication channel. Aside from message delivery, despite support from national CSIRTs, researchers in (Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]; Van Hove et al., 2023) still encounter the situation that certain stakeholders did not comply with the CVD as best practice in the first place. The feedback from stakeholders is also not always friendly, either in public or private discussions (Vanhoef, [n. d.]; Moura and Heidemann, 2023; Van Hove et al., 2023). In certain cases, to perform timely disclosure to stakeholders, researchers still need to put in extra efforts in contacting the vendors directly despite having support organizations (Van Hove et al., 2023).
@@
Furthermore, reviewing and responding to stakeholders can take time and effort for the finders. With the ticketing systems as the common practices from stakeholders, the automatic responses in large-scale notifications can result in a high amount of message content to be examined (Çetin et al., 2017a; Li et al., 2016; Stock et al., 2018b, 2016a; Chen et al., 2024). Even though automatic messages may share patterns to be categorized, the mixture of multiple languages(Li et al., 2016; Boucher and Anderson, 2022; Chen et al., 2024; Çetin et al., 2017a), unclear stakeholder disclosure policies (Chen et al., 2024; Stivala et al., 2024), communication systems requiring manually efforts to register and input messages (Çetin et al., 2017a; Chen et al., 2024) may bring higher than expected workload to finders. This can hinder the effectiveness of disclosure and notification, and more importantly, the incentives of finders (Çetin et al., 2017a; Boucher and Anderson, 2022; Moura and Heidemann, 2023; Chen et al., 2024).
6. Best Practices for Disclosure and Notification
@@
In this section, we aim to examine the gap in best practices and propose a new best practice for large-scale vulnerability disclosure and notification based on what we have learned from selected work, communities, and support organisations. We follow the same structure as in the previous section to look into limitations and opportunities in the stages of disclosure and notification operations. In each stage, we discuss the common pitfalls to avoid, tradeoffs on method selection, and eventually, the suggestions to both a finder and other stakeholders on how to adopt the best practice at scale.
@@ 6.1. Pre-Assessment
Understanding the impact scale, vulnerability characteristics, vulnerability disclosure or notification, and potential stakeholders is essential to setting up a disclosure operation. Whether it’s vulnerability disclosure or vulnerability notification to vendors or to end-users will lead to different communication channels, messaging infrastructure, disclosure policies, and message content. From what we observed and as revealed in the reflections of selected work, due to the lack of large-scale vulnerability disclosure and notification best practice, a finder may not fully understand efforts and tradeoffs to conduct the disclosure during the pre-assessment stage (Boucher and Anderson, 2022; Moura and Heidemann, 2023). For instance, researchers, ethical hackers, and practitioners may not be aware of the difference between disclosure to vendors or notification to end-users, prioritisation of the contact list, and the coverage of national CSIRT support. They then may face stress or even frustration once they start the disclosure operations, then struggle with unexpected challenges (Moura and Heidemann, 2023). Therefore, timely consultation with CSIRTs or equivalent institutes can help a finder comprehend the potential impact of vulnerability and the scale of the notification to stakeholders, as revealed in the timeline in (Van Hove et al., 2023) and suggested by several support organisations (Forum of Incident Response and Security Teams (2020), FIRST; oec, 2021; National Cyber Security Centre of The Netherlands (2018), NCSC-NL; CERT/CC, 2024). Moreover, this can also help a finder evaluate potential stakeholder responses. Finders and receiving parties may have different risk assessment standards, resulting in different definitions of vulnerability severity from both ends (sam, 2024; mic, 2016). As revealed in reviewed work (Li et al., 2016; Durumeric et al., 2014), other stakeholders may set a lower priority or don’t have enough capacity to mitigate the issues earlier than the finders expect. Finally, the next common question is which combination of channel, messaging infrastructure, disclosure policy and interaction can be the most effective in each case. These aspects will be discussed in the following stages.

### v2

@@ 4.2.3. Messaging Infrastructure
A finder will build their messaging infrastructure or delegate it to support organisations based on the selected communication channel. The common practice of outgoing email and website is using a registered individual or disclosure-specific account with the domain name of the finder’s organisation to increase the delivery rate and trust of recipients, as implemented in (Li et al., 2016; Stock et al., 2016a, 2018b; Chen et al., 2024). The email account selection of each selected publication is presented in Table 3. The dedicated account practice is to address the email notification challenge of spam and phishing mitigation. Due to the increasing spam and phishing messages, more and more measures are in place to prevent the delivery of these messages. In the early 2000s, it was possible to spin up a mail server and immediately send out an email notification to thousands of recipients. These days, there are many different standards associated with sending out emails (SPF, DKIM, DMARC, etc.), as mentioned in (Li et al., 2016), and the reputation of the mail server is taken into account before an email is delivered.
@@
Although a finder or its organisation can maintain their messaging infrastructure, it is also reported in (Chen et al., 2024) that the finders’ email infrastructure is partially outsourced to a mail server, Microsoft Exchange. Their mail server introduces several restrictions, such as limited account control and email-sending rate, which prevent sending out large amounts of messages (Microsoft, [n. d.]) efficiently. This makes it harder for finders to perform large-scale notification operations using email. This also resonates with the trends of vendors outsourcing vulnerability platforms reported in (Boucher and Anderson, 2022), which limit the reachability from the sender to the responsible parties, the client of the outsourced platforms.
@@
There are disclosure and notification operations delegating the messaging to national CSIRTs (such as NCSC-NL), CERT/CC with VINCE (CERT/CC, [n. d.]), and The Shadowserver Foundation. These are listed as Coordinator Delegation and The Shadowserver Foundation in the Table 3. While VINCE serves as a disclosure database with notification to vendors, as adopted in (Boucher and Anderson, 2022; Vanhoef and Piessens, 2017a), national CSIRTs have their mailing list or website to inform vendors and end-users, which was mentioned in (Li et al., 2016; Stock et al., 2016a; Van Hove et al., 2023; Beitis and Vanhoef, [n. d.]). In particular, The Shadowserver Foundation has become a newer notification delegation option with the organisation’s own messaging infrastructure to inform the stakeholders, including vendors and end-users, such as ISPs, domain owners, and network operators (Sha, [n. d.]). Besides, if a finder works with an organisation with dedicated communication channels, such as an ISP notifying its customers with its internal notification and management tool in (Çetin et al., 2019b; Rodríguez et al., 2021), there can be a more efficient way to trigger action from the recipients. Furthermore, finders in (Li et al., 2016; Çetin et al., 2017a; Chen et al., 2024) observed that stakeholders, such as domain name owners and cloud providers, used their notification systems to prompt their customers to take action.
4.2.4. Disclosure Policy and Message
@@
The formulation of the disclosure message can determine the attention of the recipients and whether or not to respond and take action (Wogalter et al., 1999). The tradeoffs of length and details of content, such as remediation and security suggestions, are extensively discussed in (Çetin et al., 2017a; Stock et al., 2016a, 2018b; Çetin et al., 2019b; Li et al., 2016; Chen et al., 2024; Stivala et al., 2024), where the authors document not only their disclosure message templates but also the feedback from stakeholders.
@@
Finders in (Çetin et al., 2017a; Stock et al., 2016a, 2018b) provided brief vulnerability information, affected systems, and disclosure purposes in their initial messages. The messages prompted recipients to either respond to the emails or visit a webpage for more information on vulnerabilities and remediations, using the automatically generated token from each message provided by the finders. This way, the finders could monitor the response rate with a dedicated web backend. Moreover, the finders can reduce the risk of information leaks in cases involving incorrect recipients. Furthermore, disclosure messages could be embedded with HTML content, such as the logo of the finder’s organisations, to further check if a recipient loads the full message. However, due to the spam filter and recent email client loading mechanism, extensive embedded HTML content is no longer a suggested method confirmed in (Stock et al., 2018b; Çetin et al., 2017a). From the common experiences of earlier (Çetin et al., 2017a; Stock et al., 2018b; Li et al., 2016) to recent publications (Stivala et al., 2024; Chen et al., 2024), the finders confirmed that plain text is the suggested way with less distraction and distinction from phishing messages.
@@
In large-scale vulnerability notifications, providing information such as domain names, IP addresses, ports, and issues found in the vulnerable systems is recommended to help stakeholders investigate issues in time, as documented from the stakeholders’ feedback in (Li et al., 2016; Stock et al., 2018b; Çetin et al., 2017a; Chen et al., 2024). However, receiving parties may have certain mail filters for incoming messages, which strictly limit the message length and attachments, as reported in (Chen et al., 2024). This could hinder the case of cloud providers or domain owners with a large number of vulnerable systems running for their clients. Such stakeholders may decline the messages containing longer vulnerable host lists or attachments. This eventually makes the decision on the message content more challenging for finders (Çetin et al., 2017a; Chen et al., 2024; Stivala et al., 2024).
@@
Organisations present support for CVD by publishing a disclosure policy on their websites, using security.txt(Foudil and Shafranovich, 2022), or providing bounty or tracking programs (Boucher and Anderson, 2022; Moura and Heidemann, 2023; Vanhoef, [n. d.]). Boucher and Anderson (Boucher and Anderson, 2022) observed vendors outsourcing their programs to third-party platforms that reveal their own policies and prioritisation on vulnerability selections. They noted that the policies and preferences of the outsourced platforms may limit the incentive and direct message delivery to the responsible parties. In the vulnerability notification, Chen et al. (Chen et al., 2024) observed that most of the stakeholders still do not provide security.txt or equivalent information in their responses, but privacy policies that do not necessarily help the disclosure process. The implementation of the disclosure policies among stakeholders still requires attention.
@@
Google Project Zero (Pro, [n. d.]) started with an outgoing vulnerability disclosure policy describing the timelines they would use in disclosing vulnerabilities. On the one hand, this has pressured vendors to work on mitigation and publish it within 90 days of the timeline. After several years, the 90-day deadline has become an accepted practice. Academic researchers have also started using outgoing vulnerability disclosure policies (Reidsma et al., 2023), which help build trust between stakeholders and coordinate disclosure operations. The outgoing policy has been implemented in (Van Hove et al., 2023; Chen et al., 2024), where the finders presented their intended procedures on vulnerability handling, notification frequency, public disclosure schedule, and more (National Cyber Security Centre of The Netherlands (2018), NCSC-NL; van der Ham et al., 2023) in their outgoing messages. This gives the receiving party a brief yet informative message for potential procedures and conversation. Other finders in (Çetin et al., 2017a, 2019b; Stock et al., 2018b; Moura and Heidemann, 2023; Stivala et al., 2024) did not document a dedicated disclosure policy, yet did provide equivalent information and contact methods in their messages for coordinating or exempting from the notifications. As for the finders that chose coordinator delegation, ISP intervention, and The Shadowserver Foundation, they also followed and presented the disclosure policy from the support organizations as documented in (Boucher and Anderson, 2022; Çetin et al., 2019b; Van Hove et al., 2023; Moura and Heidemann, 2023; Rodríguez et al., 2021; Stock et al., 2018b, 2016a; Vanhoef and Piessens, 2017a; Vanhoef, [n. d.]). These delegated parties often use existing relations with stakeholders and customers to create a trustworthy communication channel.
@@ 4.2.5. Post-Review
The remediation rate of the vulnerable systems is a gripe in most selected publications, as shown in Table 2. It is worth noting that the remediation rates of each selected publication do not directly indicate the efforts and success of the channel selection and messaging infrastructure from the finders. Severity of reported vulnerabilities, impact scale, affected system, and risk management of receiving parties will all contribute to the stakeholders’ actions and mitigation progress. Although 57% of the vulnerable hosts patched with email notification in (Durumeric et al., 2014), which is relatively higher than other notification operations with remediation rates from lower than 2.25% to up to around 30% (Çetin et al., 2017a; Stock et al., 2018b, 2016a; Li et al., 2016; Chen et al., 2024; Stivala et al., 2024). Durumeric et al. (Durumeric et al., 2014) initiated the notification of the Heartbleed vulnerability 3 weeks after the notable public disclosure. Besides, they had to drop 56% of the detected vulnerable hosts due to responsible administrators likely having no access to treat the embedded devices (Durumeric et al., 2014). On the contrary, with a dedicated communication channel and trustworthy notification organisation (ISP Intervention) as in (Çetin et al., 2019b; Rodríguez et al., 2021), the remediation rate can be significantly higher for more than 75%; nonetheless, such a scenario requires selective recipients or extra capacity and won’t necessarily fit other end-user notifications (Çetin et al., 2019b). Moreover, not every disclosure operation can measure the remediation rate due to the nature of the vulnerabilities and affected parties (Boucher and Anderson, 2022; Van Hove et al., 2023; Moura and Heidemann, 2023; Vanhoef and Piessens, 2017a). However, Beitis and Vanhoef (Beitis and Vanhoef, [n. d.]) include both disclosure and notification operations with a measurable impact scale. They did not provide a remediation rate since their work was still a work in progress. Their result is worth observing in the near future.
@@
The lack of best practices in large-scale vulnerability disclosure, particularly in notification, among academic researchers and practitioners, has led to the aforementioned struggles and challenges at each stage, as noted in the title of (Moura and Heidemann, 2023), ”Vulnerability Disclosure Considered Stressful”. The practice gaps motivated the finders in the selected publications to adopt existing guidelines, reflect on real operations, and contribute to best practices. However, the stress and frustration of the finders deserve attention. Whether it is large-scale disclosure or notification, various vulnerability platforms (Boucher and Anderson, 2022), tracking systems (Moura and Heidemann, 2023), and ticketing systems (Li et al., 2016; Çetin et al., 2017a; Chen et al., 2024) have increased the workload of finders to deliver the messages to the responsible parties in the complex multi-party scenarios. The inaccuracy of existing abuse and generic contact information has caused false positives in contact retrieval and information leaks to unintended recipients, which is mentioned in nearly all selected publications using email as a communication channel. Besides, despite support from national CSIRTs, finders in (Vanhoef and Piessens, 2017a; Van Hove et al., 2023) still encountered the situation that certain stakeholders did not comply with the CVD as best practice in the first place. The feedback from stakeholders is also not always friendly, either in public or private discussions, as reported in (Vanhoef, [n. d.]; Moura and Heidemann, 2023; Van Hove et al., 2023). In certain cases, to perform timely disclosure to stakeholders, finders still need to put in extra efforts in contacting the vendors directly despite having support organisations (Van Hove et al., 2023) or disclosure programs presented by stakeholders (Boucher and Anderson, 2022).
@@
Furthermore, reviewing and responding to stakeholders can take time and effort for the finders. With the ticketing systems as the common practices from stakeholders, the automatic responses in large-scale notifications can result in a high amount of message content to be examined, which is documented in (Çetin et al., 2017a; Li et al., 2016; Stock et al., 2018b, 2016a; Chen et al., 2024). Even though automatic messages may share patterns to be categorised, the mixture of multiple languages in messages (Li et al., 2016; Boucher and Anderson, 2022; Chen et al., 2024; Çetin et al., 2017a), unclear stakeholder disclosure policies (Chen et al., 2024; Stivala et al., 2024), and stakeholders’ communication systems requiring manual efforts to register and input messages (Çetin et al., 2017a; Chen et al., 2024) may bring higher than expected workload to finders. This can hinder the effectiveness of large-scale disclosure and notification, and more importantly, the incentives of finders, as discussed in (Çetin et al., 2017a; Boucher and Anderson, 2022; Moura and Heidemann, 2023; Chen et al., 2024).
To sum up, we conclude by drawing insights from our operations stage model, which is extracted from the selected publications and the experiences shared by support organisations and other communities. This also fulfils our third research question presented in Section 1.1. Furthermore, the insights have highlighted current gaps in best practices for large-scale vulnerability disclosure, particularly in notification, across both the academic security research community and other communities.
5. Best Practices for Large-Scale Vulnerability Disclosure and Notification
@@
In this section, we aim to examine the gap in current guidelines and propose new best practices for large-scale vulnerability disclosure and notification based on what we have learned from selected academic publications, communities, and support organisations. The main focus is to help the finders from communities that include academic security researchers, practitioners, and ethical hackers. Yet, the best practices are not limited to finders but also stakeholders to improve the receiving strategies to handle vulnerability reports. We follow the same structure as in the previous section to look into limitations and opportunities in the stages of disclosure and notification operations. In each stage, we discuss common pitfalls to avoid, tradeoffs in method selection, and provide suggestions to the finder and other stakeholders on adopting our best practices.
@@ 5.1. Pre-Assessment
Understanding the impact scale, vulnerability characteristics, vulnerability disclosure or notification, and potential stakeholders is essential for a finder to set up a disclosure or notification operation. Whether it’s vulnerability disclosure or vulnerability notification to vendors or to end-users will lead to different communication channels, messaging infrastructure, disclosure policies, and message content. From what we observed and as revealed in the reflections of selected publications, due to the lack of large-scale vulnerability disclosure and notification best practices for different communities, a finder may not fully understand efforts and tradeoffs to conduct the disclosure during the pre-assessment stage, which are documented in (Boucher and Anderson, 2022; Moura and Heidemann, 2023). We list the points below to highlight the common pitfalls and our suggestions:
  • Security researchers, ethical hackers, and practitioners may not be aware of the difference between disclosure to vendors or notification to end-users, prioritisation of the contact list, and the coverage of national CSIRT support. They may then face stress or frustration during the disclosure operations and struggle with unexpected challenges, as noted by Moura and Heidemann (Moura and Heidemann, 2023).
  • A finder and a receiving stakeholder may have different risk assessment standards, resulting in different definitions of vulnerability severity from both ends (sam, 2024; mic, 2016). As revealed in (Li et al., 2016; Durumeric et al., 2014), the receiving parties may set a lower priority or lack the capacity to mitigate the issues earlier than the finders expect.
  • We recommend timely consultation with CSIRTs or equivalent support organisations. This can help a finder comprehend the potential impact of vulnerability and the scale of the notification to stakeholders, as revealed in the timeline by van Hove et al. (Van Hove et al., 2023) and suggested by support organisations (Forum of Incident Response and Security Teams (2020), FIRST; oec, 2021; National Cyber Security Centre of The Netherlands (2018), NCSC-NL; CERT/CC, 2024; BOD, 2020). Moreover, this can also help a finder estimate potential stakeholder responses and prepare the message handling in advance.
The next common question is which combination of channel, messaging infrastructure, disclosure policy and interaction can be the most effective in each case. These aspects will be discussed in the following stages.